OpenStudy (anonymous):

plz plz plz help! find the limit. lim x--->2 4-(radical 18-x)/(x-2)

OpenStudy (anonymous):

\[\large lim_{x\rightarrow2}\frac {4-\sqrt{18-x}}{x-2} \] this the question?

OpenStudy (anonymous):

if yes... you'll need to rationalize the NUMERATOR....

OpenStudy (anonymous):

okay so its 4-(Sq.rt 18-x)/x-2 and then ill multipl. by -4+(sq.rt 18-x/x-2 to the top and bottom right?

OpenStudy (anonymous):

you'll multiply this to the top and bottom: \[\large 4+\sqrt{18-x} \]

OpenStudy (anonymous):

yeah and i'll get -16-18+x/-4x+18-x+8-2(sq rt 18-x/ right?????

OpenStudy (anonymous):

i mean sq rt 18-x) right?

OpenStudy (anonymous):

can you give it to me simplified?

OpenStudy (anonymous):

ummm okay 34+x/-4x+(sq rt 18-x)

OpenStudy (anonymous):

i mean -34

OpenStudy (anonymous):

hmmm.. i got something different... lemme show what i got...

OpenStudy (anonymous):